OpenStudy (anonymous):

1. (2r^2-3r-35) / (r-5) 2. (3p^2+20p+11) / (p+6) 3. (n^2+7n+12) / (n+3) #4 will be drawn

OpenStudy (anonymous):



OpenStudy (luigi0210):

Use synthetic division

OpenStudy (anonymous):

whats that

OpenStudy (anonymous):

i only know FOIL method

OpenStudy (anonymous):

and theres another one i forgot it

OpenStudy (luigi0210):

Synthetic is upside down division

OpenStudy (anonymous):

oh i just looked up what synthetic division is on google. we don't use that.

OpenStudy (luigi0210):

Well trust me, that is the easiest way

OpenStudy (anonymous):

@Luigi0210 there is another way it is like this(look at drawing) i know how to start it but not finish it:

OpenStudy (anonymous):



OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (goldphenoix):

\[\frac{ (2r^2-3r-35) }{ (r-5) }\] \[\frac{ (3p^2+20p+11 }{ (p+6) }\] \[\frac{ (n^2+7n+12 }{ (n+3) }\] Gonna make it as a reference ( Make life easier)

OpenStudy (anonymous):

what do u mean reference? and how do u do that fancy txt? o.o

OpenStudy (anonymous):

and u forgot #4. it is drawn up there^ if u didnt see, just saying

OpenStudy (goldphenoix):

Reference as in make your problem easier to understand. Fancy txt = Equation button underneath.

OpenStudy (goldphenoix):

LOL. My bad.